# references.bib

### v2

@@ 3.1.1 Design of a simple kit for collecting trichomes from tomato leaves
In this study, we designed a straightforward kit to measure the density of type VI glandular trichomes on tomato leaves. This kit comprises custom-made measurement paper crafted from inkjet-printable cardstock and cellophane tape (NICHIBAN, CT-18), enabling cost-effective measurements (Figure 4). This paper contains augmented reality (AR) markers, enabling the distance of objects within captured images to be precisely measured. Additionally, there is a 12 mm $\times$ 12 mm opening in the center of the paper. A piece of cellophane tape is affixed to the back of the paper, covering this opening. This tape is used to extract trichomes from the surface of tomato leaves.
@@
Optimizing a technique that involves the use of cellophane tape for trichome extraction allows the creation of a highly cost-efficient kit. Although recent advances in sophisticated object detection algorithms have utilized complex neural networks, these neural networks typically demand high-speed parallel processing, considerably increasing computational costs. Our approach, which uses cellophane tape for trichome extraction, circumvents the need for such intensive computational resources and enables the use of less computationally demanding algorithms for trichome detection. Thus, this approach reduces the overall cost of the assessment process, making it more economically viable.
The purpose of this diagnostic kit is to meet the technological progress needs in agricultural production areas that lack access to laboratory-level experimental facilities. Therefore, to minimize the number of components required for measurement, the kit is specifically designed to allow for dual manual operation. The measuring paper is held in one hand, whereas the smartphone is operated in the other hand (Figure 4). The statistical model accounts for the variability in the results caused by varying optical conditions among shots. The kit is designed to prioritize user friendliness as a minimum viable product in lean startup methodology for manufacturing and software development, enabling the measurement of trichome density without requiring technical expertise [mu2013lean, Ries2011-xn]. This technology is well suited for practical application in the field because of its simplicity and cost-effectiveness (Figure 5).
@@ 3.1.2 Proposed concept of the tomato cultivation support framework
This kit is designed to empower farmers to assess the additional fertilization needs of their tomato plants based on trichome density and receive fertilization recommendations. The proposed framework involves a four-step process. First, trichomes are sampled from tomato leaves via adhesive cellophane tape on measurement paper. Second, a smartphone is used to capture an image of the measurement paper with the transferred trichomes. Third, the captured image is analyzed to determine trichome density, which serves as an indicator of the degree of fertilizer stress. Finally, on the basis of the analyzed trichome density, the system provides a fertilization recommendation to the farmer, guiding their decision on whether additional fertilization is necessary.
@@ 3.1.3 Overview of the trichome detection image processing pipeline
This study implemented an image processing pipeline for calculating trichome density. Instead of using deep learning models for direct image processing, tabular data analysis was chosen. Although deep learning often requires substantial computational resources, this architecture leverages the inherent simplicity and reduced dimensionality of tabular data, enabling the use of computationally lighter models. This method effectively minimizes computational costs while bolstering quantitative model interpretation, thereby improving the reliability and interpretability of the results for agricultural applications.
To avoid data leakage, an image processing pipeline was developed using images obtained from preliminary experiments. This approach ensured that the model development and evaluation processes remained independent and unbiased.
The image processing pipeline comprises three main stages. The preprocessing stage involves extracting metadata in the exchangeable image file format (EXIF) embedded in the image files, identifying AR markers within the images, applying homography transformations for image alignment, and adjusting for variations in lighting conditions. This preprocessing stage also includes binarizing images to segment trichomes and applying noise removal filtering algorithms to enhance trichome visibility.
In the feature extraction stage, we focused on identifying trichomes in the images, specifically targeting type VI glandular trichomes, owing to their unique morphology. Furthermore, AR markers were utilized to convert pixel counts to relative distances, enabling accurate distance estimation. The final stage, density calculation, involves calculating the trichome density from images captured under various environmental conditions and camera settings, each associated with distinct EXIF data. Algorithm 1 presents an outlined function of the trichome density computation process using our custom-developed measurement kit. The image processing procedures described above were executed via the software specifications outlined in Table 3.
@@ 3.1.4 Image preprocessing: AR marker detection, illumination correction, and noise reduction
The first step of the developed image processing pipeline is to read EXIF data from the image files. This metadata contains valuable information about the environmental conditions and camera settings at the time of image capture. By analyzing aspects such as exposure time and ISO speed ratings, we can gain a comprehensive understanding of the variations in imaging conditions and their potential impacts on trichome density measurements.
To measure the absolute scale and orientation of the captured images, we incorporated AR markers into our image processing pipeline. These markers enabled us to determine the trichome density of the captured images. For AR marker detection, we utilized the ArUco module from the OpenCV library (version 4.6.0.66) [intelcorporationopencvcontribpython].
The AR marker detection algorithm employs a simple, nonrecursive approach, as the development of AR technology was not the primary focus of this study. The process involves image preprocessing, where the captured image is converted to grayscale and then binarized via Otsu’s method to enhance marker visibility under various lighting conditions. We used a predefined "ArUco dictionary", DICT_4X4_250. This dictionary offers a balance between the number of unique markers and detection robustness. Finally, the OpenCV function cv2.aruco.detectMarkers() is used to identify the ArUco markers in the preprocessed image.
The complete algorithm is presented in Algorithm 2.
@@
Following AR marker detection, we apply homography transformation to correct perspective distortion in images. This distortion is caused by positional and orientational variations in the measurement paper relative to each camera captured. The homography transformation adjusts the images as though the camera and the measurement paper were parallel. This process involves using the coordinates of the four corners of a region on the measurement paper, obtained from detecting AR markers as reference points to estimate a perspective transformation matrix. By applying this transformation, we ensure that the images are standardized for accurate trichome density analysis, regardless of slight variations in camera positioning during image capture.
After applying the homography transformation, we implemented an illumination correction procedure to address nonuniform lighting conditions across the acquired images. This step is essential for maintaining consistency in trichome detection, regardless of ambient lighting variations during image acquisition.
Our approach uses a flat field correction technique [seibert1998medical], which involves several preprocessing steps. Initially, all the images are resized to a standardized dimension of 1000 pixel $\times$ 1000 pixel to ensure uniform processing. We subsequently apply an averaging filter with a kernel size of 50 pixel $\times$ 50 pixel to generate a smoothed version of the image, representing the overall illumination pattern.
The correction is mathematically described by the following equation:
$C=\frac{R\times m}{F}$ (1)
@@
39:$C\leftarrow\frac{R\times m}{F}$
@@ 3.1.5 Feature extraction: trichome identification and density analysis
Following noise removal, a segmentation procedure based on the watershed algorithm was implemented to identify individual trichomes. This method detects boundary surfaces that separate contours in the image. The watershed algorithm was chosen because of its suitability for rule-based segmentation on the basis of the distinctive appearance of trichomes, offering a computationally efficient solution. Although deep learning-based segmentation can often achieve high accuracy, it typically requires substantial computational resources, often necessitating the use of devices with advanced parallel processing capabilities, such as graphics processing units, to be integrated into the computer system. In contrast, the watershed algorithm, which is a traditional image-processing technique, generally operates with lower computational costs and can often perform effectively on standard computer hardware, such as central processing units. This characteristic is particularly important for our application, as it ensures that the tomato nutrient diagnostic kit remains cost-effective for farmers, aligning with our goal of developing accessible and practical tools for agriculture.
@@
In this paper, we utilize the term "trichomes" to describe the contours of detected objects based on the following criteria:
@@
  • The perimeter of the object falls within a specified range, defined by a lower and upper limit;
  • The area of the object lies within a specified range, defined by a lower and upper limit.
The contours detected through segmentation yield centroid points, denoted as in Equation 2. These points exist within a two-dimensional space and are represented via the pixel coordinates of the image. Figure 6 shows the image processing steps performed up to this point, including preprocessing and object detection.
The watershed segmentation process yields a set of detected contours, from which we extract centroid points. These points are represented in a two-dimensional space corresponding to the pixel coordinates of the image. The set of centroids is denoted as $P$, as defined in Equation 2.
$P={p_{i}\in\mathbb{R}^{2}\mid i=1,2,\ldots,N}$ (2)
where $p_{i}$ represents the i-th centroid point in the two-dimensional real space $\mathbb{R}^{2}$, and $N$ is the total number of detected contours. Each point $p_{i}$ can be expressed as $(x_{i},y_{i})$, where $x_{i}$ and $y_{i}$ are the horizontal and vertical pixel coordinates, respectively.
@@
19:     Calculate circularity, perimeter, and area for all contours
@@
The detailed process of trichome identification, including the criteria and thresholds used for filtering, is presented in Algorithm 4. This algorithm encapsulates the key steps and decision-making process involved in distinguishing type VI glandular trichomes from other objects in the image. Figure 6 shows the image processing steps performed up to this point, including preprocessing and object detection.
@@
The density was computed for the centroid points denoted as $P$ in Equation 2. We employed the nearest neighbor distance (NND) method, as described in Equation 3 [clark1954distance], to quantify the spatial distribution of these points. The NND method involves measuring the distance between each point and its nearest neighbor within the set.
$NND=\frac{1}{n}\sum_{i=1}^{n}d(p,q)$ (3)
In this equation, $d(p,q)$ represents the Euclidean distance between point $p$ and its closest neighboring point $q$. The NND is then calculated by averaging these distances over the total number of contours $n$ present in the image. Efficiently computing the nearest neighbors of each point is crucial, especially when many points are dealt with. A brute-force approach, where distances between all point pairs are calculated, would have a computational complexity of $O(nd)$, where $n$ is the number of data points and $d$ is the dimensionality. This "big O" notation, $O(nd)$, describes how the execution time of the algorithm increases as the input size ($n,d$) increases[skiena1998algorithm]. In this case, $O(nd)$ means that the execution time increases linearly with the number of points and the dimensionality. Therefore, the brute-force approach can become computationally expensive as the number of points increases. To address this issue, we utilized a k-dimensional tree algorithm, which organizes the points in a space-partitioning data structure that allows for rapid nearest neighbor queries. This algorithm effectively reduces the computational complexity to $O(d\log n)$ [cunningham2022knearest], meaning that the execution time grows linearly with the dimensionality but only logarithmically with the number of points. This logarithmic growth makes the k-d tree algorithm much more efficient than the brute-force approach, especially for large datasets. This algorithm is particularly suitable for our two-dimensional image analysis ($d=2$) with many trichome centroids. The detailed implementation of the NND calculation is presented in Algorithm 5.
@@
The NND serves as an inverse indicator of contour density, with higher values denoting sparse arrangements and lower values signifying denser groupings. Glandular type VI trichomes are susceptible to detachment from leaf surfaces because of factors such as insect infestations or physical abrasion. However, NND demonstrates superior robustness to such accidental detachment compared with conventional density indices, which are based on the object count per unit area. This robustness arises because NND represents density as a probability distribution, which enables it to maintain reliable estimates even when localized areas experience trichome detachment, provided that unaffected regions persist. Notably, the NND calculation requires the presence of at least two points. The appendix elucidates this concept through a simulation employing point cloud data generated via a Poisson process in two-dimensional space.
@@ 3.2.1 Selection of the cultivation method
A well-controlled cultivation experiment is essential for evaluating the proposed method for assessing the nutrient status of tomato plants with testability. This cultivation experiment aimed to measure phenotypes such as trichome density and leaf nitrate ion concentration and fruit yield under ideal conditions for tomato growth, minimizing the variability inherent in soil-based cultivation. Therefore, we opted for hydroponics rather than soil-based cultivation to circumvent the complexities associated with soil properties such as variations in soil composition and structure. These variations can lead to complexities in cultivation experiments, including inconsistent water retention, fertilizer dynamics, and potential for pest and disease outbreaks. This decision allowed for more precise control over nutrient availability and minimized potential confounding factors in our evaluation of fertilizer stress. This approach also enabled us to investigate the relationships between nutrient levels and plant responses more clearly in an ideal growth environment that approximates the optimal growth conditions found in some commercial cultivation systems.
Among various hydroponic techniques, we selected the deep water culture (DWC) method for its simplicity, minimal equipment requirements and ease of implementation in a randomized block design. In DWC, plant roots are fully submerged in nutrient solution within a production tank, which is aerated to sufficiently oxygenate the roots [hoagland1938the, velazquezgonzalez2022a]. This straightforward approach facilitates the relocation and replication of experimental units regardless of location, contributing to increased falsifiability. Moreover, DWC allows easy control of nutrient solution conditions at the individual plant level, making it suitable for randomized block designs.

### v3

@@ 3.1.1 Design of a simple kit for collecting trichomes from tomato leaves
To establish practical nutrient assessment methods for young tomato leaves, we developed a straightforward measurement kit that addresses the need for young leaf measurements, the limitations of existing optical methods in the presence of trichomes, and the need for field-applicable solutions. The kit was designed to measure the density of type VI glandular trichomes on tomato leaves as indicators of nutrient status.
To ensure accessibility and practical application in agricultural settings, we designed the kit using readily available materials, specifically custom-made measurement paper crafted from inkjet-printable cardstock and cellophane tape (NICHIBAN CT-18). The selection of cellophane tape was driven by the three-dimensional characteristics of tomato leaves. The leaf veins and natural leaf curling create macroscopic undulations with longer periods than those caused by trichome surface structures. Given these topographical features, we needed a measurement method compatible with cameras that have a limited depth of field. Cellophane tape, which can closely conform to the leaf surface, was selected as both a thin-film substrate and transfer medium, enabling the conversion of three-dimensional trichome distributions into measurable two-dimensional patterns. This design choice enables cost-effective measurements while maintaining reliability (Figure 4).
To achieve consistent measurements across varying field conditions, the measurement paper incorporates augmented reality (AR) markers, which enable precise distance calibration within captured images, regardless of the camera position or environmental conditions. Additionally, there is a 12 mm $\times$ 12 mm opening in the center of the paper, which is covered by the cellophane tape for trichome extraction.
@@
The development of this measurement approach was guided by the practical constraints of agricultural applications. While recent advances in object detection have predominantly relied on complex neural networks requiring high-speed parallel processing, our optimized approach using cellophane tape for trichome extraction circumvents these computational demands. This technique, combined with simplified image processing algorithms, creates a practical solution that maintains accuracy while eliminating the need for intensive computational resources, which makes the assessment process more economically viable for agricultural applications.
This diagnostic kit was developed to meet technological needs in agricultural production areas with limited access to laboratory facilities, fulfilling the requirement for reliable nutrient assessment tools in field conditions. To minimize the number of components required for measurement, the design enables dual manual operation, with the measuring paper in one hand and the smartphone in the other (Figure 4), while maintaining measurement reliability through statistical modeling that accounts for varying optical conditions. In accordance with lean startup methodology principles [mu2013lean, Ries2011-xn], the kit was developed as a minimum viable product that prioritizes user-friendly operation without requiring technical expertise, enabling rapid assessment of plant nutrient status in the field (Figure 5).
@@ 3.1.2 Proposed concept of the tomato cultivation support framework
To translate trichome density measurements into practical agricultural decisions, we developed a systematic support framework that empowers farmers to assess the nutritional needs of their tomato plants. This framework streamlines the assessment process through four sequential steps. The process begins with sampling trichomes from tomato leaves via adhesive cellophane tape on measurement paper. Next, a smartphone is used to capture an image of the transferred trichomes. The captured image is then analyzed to determine trichome density as an indicator of fertilizer stress. Finally, specific fertilization recommendations are generated based on the analysis results. This structured approach enables farmers to make informed decisions about whether additional fertilization is necessary, replacing subjective assessments with objective evaluations.
@@ 3.1.3 Overview of the trichome detection image processing pipeline
The success of this support framework critically depends on accurate and reliable trichome density measurements. To realize such measurements while maintaining practical utility in agricultural settings, we developed an image processing pipeline that prioritizes computational efficiency and result interpretability. Rather than employing deep learning models for direct image processing, we chose a hybrid approach that combines traditional computer vision techniques with tabular data analysis. This architecture leverages the inherent simplicity and reduced dimensionality of tabular data, effectively minimizing computational costs while bolstering quantitative model interpretation. This approach enables the use of computationally lighter models while improving both the reliability and interpretability of the results for agricultural applications.
Our development process prioritized methodological rigor and practical reliability. To avoid data leakage, we developed the image processing pipeline using images from preliminary experiments, ensuring that the model development and evaluation processes remained independent and unbiased.
The pipeline consists of three main stages, beginning with preprocessing, where we extract metadata in the exchangeable image file format (EXIF) embedded in the image files. At this stage, we identify AR markers within the images for initial alignment, apply homography transformations, and adjust for variations in lighting conditions. Additional preprocessing steps include image binarization for trichome segmentation and noise removal filtering to enhance trichome visibility.
In the feature extraction stage, we focus on identifying trichomes in the images, specifically targeting type VI glandular trichomes owing to their unique morphology. Using the aligned images and calibrated AR markers from the preprocessing stage, we convert the pixel counts to relative distances, enabling accurate distance estimation. The final density calculation stage leverages the extracted features to determine trichome density from images captured under various environmental conditions and camera settings, each associated with distinct EXIF data. The computational process is detailed in Algorithm 1, with the software specifications provided in Table 3.
@@ 3.1.4 Image preprocessing: AR marker detection, illumination correction, and noise reduction
Image processing begins with reading EXIF metadata from image files. This processing provides crucial information about environmental conditions and camera settings during image capture. Analysis of specific parameters, such as exposure time and ISO speed ratings, enables us to account for variations in imaging conditions that affect trichome density measurements, ensuring consistency across different environments.
Our spatial calibration incorporates AR markers into the image processing pipeline. These markers serve as reference points for determining the scale and orientation of the captured images, enabling trichome density determination. For marker detection, we implemented a nonrecursive approach via the ArUco module from the OpenCV library (version 4.6.0.66) [intelcorporationopencvcontribpython], as the development of AR technology was not the primary focus of this study.
For robust detection under field conditions, we implemented a systematic image preprocessing approach. The algorithm converts the captured image to grayscale, followed by binarization via Otsu’s method to increase marker visibility across lighting conditions. We selected the "ArUco dictionary" DICT_4X4_250 due to its balance between marker uniqueness and detection reliability. The markers are identified via the OpenCV function cv2.aruco.detectMarkers(). The complete marker detection algorithm is presented in Algorithm 2.
@@
Following AR marker detection, we apply homography transformation to correct perspective distortion in images. This distortion arises from positional and orientational variations in the measurement paper relative to each camera captured. The transformation adjusts images as though the camera and measurement paper were parallel, using the coordinates of four corners obtained from AR markers as reference points to estimate a perspective transformation matrix. This standardization ensures accurate trichome density analysis, regardless of variations in camera positioning during image capture.
Image quality optimization begins with illumination correction, which is an essential step to address nonuniform lighting conditions across the acquired images and to maintain consistent trichome detection despite ambient lighting variations. For this purpose, we implemented a flat field correction technique [seibert1998medical]. The process starts by resizing all images to a standardized dimension of 1000 pixels $\times$ 1000 pixels, followed by applying an averaging filter with a kernel size of 50 pixels $\times$ 50 pixels to generate a smoothed version representing the overall illumination pattern. The correction is mathematically described in Equation 1.
$C=\frac{R\cdot m}{F}$ (1)
@@
39:$C\leftarrow\frac{R\cdot m}{F}$
@@ 3.1.5 Feature extraction: Trichome identification and density analysis
For individual trichome identification, we implemented a segmentation procedure based on the watershed algorithm following the noise removal step. This method detects boundary surfaces that separate contours in the image, offering an efficient approach for rule-based segmentation suited to the distinctive appearance of trichomes. While deep learning-based segmentation can achieve high accuracy, it typically requires substantial computational resources, including graphics processing units with advanced parallel processing capabilities, to be integrated into the computer system. In contrast, the watershed algorithm, which constitutes a traditional image-processing technique, operates with lower computational costs on standard computer hardware such as central processing units, making it particularly suitable for achieving our goal of developing accessible and practical tools for agriculture.
@@
In this paper, we utilize the term "trichomes" to describe the contours of detected objects on the basis of their specific properties. These criteria for identification include the following:
@@
  • The perimeter of the object falls within a specified range, defined by a lower limit and an upper limit;
  • The area of the object lies within a specified range, defined by a lower limit and an upper limit.
The contours detected through segmentation yield centroid points in two-dimensional space, represented via the pixel coordinates of the image. Figure 6 shows the image processing steps performed thus far, including preprocessing and object detection, leading to these contour detections. The set of centroids, denoted as $P$, is expressed in Equation 2.
$P=\{\bm{p}_{i}\in\mathbb{Z}^{2}\mid i=1,2,\ldots,N\}$ (2)
where $\bm{p}_{i}$ represents the $i$-th centroid point in the two-dimensional integer space $\mathbb{Z}^{2}$ and $N$ is the total number of detected contours. Each point $\bm{p}_{i}$ is represented by its pixel coordinates $(x_{i},y_{i})$, where $x_{i}$ and $y_{i}$ correspond to the horizontal and vertical positions in the image, respectively.
@@
19:     Calculate the circularity, perimeter, and area for all contours
@@
The detailed process of trichome identification, including the criteria and thresholds used for filtering, is presented in Algorithm 4. This algorithm encapsulates the key steps and decision-making process for distinguishing type VI glandular trichomes from other objects.
@@
To quantify the spatial distribution of the identified trichomes, we computed the density of the centroid points (denoted as $P$ in Equation 2) via the nearest neighbor distance (NND) method [clark1954distance]. This approach measures the distance between each point and its nearest neighbor within the set, as described in Equation 3. In this equation, $\mathrm{d}(\bm{p}_{i},\bm{q}_{i})$ represents the Euclidean distance between point $\bm{p}_{i}$ and its closest neighboring point $\bm{q}_{i}$, with the NND calculated by averaging these distances over the total number of contours $N$ present in the image.
$\mathrm{NND}=\frac{1}{N}\sum_{i=1}^{N}\mathrm{d}(\bm{p}_{i},\bm{q}_{i})$ (3)
The efficient computation of nearest neighbors becomes crucial when dealing with numerous points. A brute-force approach, which calculates the distances between all point pairs, has a computational complexity of $O(dn^{2})$, where $n$ is the number of data points and $d$ is the dimensionality. As described by this "big O" notation [skiena1998algorithm], the execution time of the algorithm increases as the input size ($n,d$) increases, making the brute-force approach computationally expensive as the number of points increases. To address this computational challenge, we adopted a k-dimensional (k-d) tree algorithm, which organizes points in a space-partitioning data structure for rapid nearest neighbor queries. This approach effectively reduces the computational complexity to $O(n\cdot d\log n)$ [cunningham2022knearest], where the execution time increases linearly with dimensionality and increases more moderately with the number of points compared to the brute-force approach. This logarithmic growth characteristic makes the k-d tree algorithm much more efficient than the brute-force approach, especially for large datasets. The algorithm is particularly suitable for our two-dimensional image analysis ($d=2$) with numerous trichome centroids. The detailed implementation of the NND calculation is presented in Algorithm 5.
@@
The NND serves as an inverse indicator of spatial density, with higher values denoting sparse arrangements and lower values representing denser groupings. While conventional density measurements simply count objects per unit area, this measure proves particularly valuable for analyzing type VI glandular trichomes, which are susceptible to detachment from leaf surfaces due to factors such as insect infestations or physical abrasion. Compared with conventional counting methods, the NND method demonstrates superior robustness to accidental detachment. The reason for this robustness is that the NND method represents density as a probability distribution, enabling reliable estimates even when localized areas experience trichome detachment, provided that unaffected regions persist. As a prerequisite for NND calculations, at least two points must be present in the measurement area. The appendix further explains this concept through a simulation employing point cloud data generated via a Poisson process in two-dimensional space.
@@ 3.2.1 Selection of the cultivation method
A well-controlled cultivation experiment is essential for evaluating our proposed method for assessing tomato plant nutrient status while maintaining testability. The aim of the experiment was to measure key phenotypes, i.e., trichome density, leaf nitrate ion concentration, and fruit yield, under ideal growth conditions. We chose hydroponics over soil-based cultivation to avoid the complexities inherent in soil properties, such as variations in composition and structure that can lead to inconsistent water retention, unpredictable fertilizer dynamics, and increased risk of pest and disease outbreaks. This approach enabled precise control over nutrient availability, and it minimized confounding factors in our evaluation of fertilizer stress while providing an ideal growth environment that approximates the optimal conditions found in commercial cultivation systems.
From the available hydroponic techniques, we selected the deep water culture (DWC) method, where plant roots are fully submerged in nutrient solution within a production tank, which is aerated to sufficiently oxygenate the roots [hoagland1938the, velazquezgonzalez2022a]. The precise control of nutrient conditions at the individual plant level makes this method particularly suitable for implementation in a randomized block design. Additionally, its simplicity of setup and minimal equipment requirements facilitate both the relocation and replication of experimental units across locations, enhancing the experiment’s falsifiability.